find the x- and y- intercepts
y = 2x + 10
finding the y
by letting x = 0
y = 2x + 10
y = 2(0) + 10
y = 0 + 10
y = 10,0
but with finding the x
by letting y = 0
y = 2x + 10
0 = 2x + 10
and I'm stuck

Found 2 solutions by checkley71, tutorcecilia:
Answer by checkley71(8403)

0=2x+10 or 2x=-10 or x=-10/2 or x=-5 or (-5,0)

Answer by tutorcecilia(2152)

but with finding the x
by letting y = 0
y = 2x + 10
0 = 2x + 10
0-10 = 2x+10-10 [Solve for x by first subtracting 10 from both sides]
-10 = 2x [Divide both sides by 2]
-10/2=2x/2 [Reduce]
-5=x
(-5,0)
So, the interceptions are (-5, 0) and (0, 10)